# spacing=nonfrench

### v2

@@ 3.2 Limitations of state-of-the-art approaches
Leveraging the PLMR model, we analyze why existing AI systems fail to fully utilize wafer-scale accelerators. To run an LLM model on a wafer-scale accelerator, we generally have two choices: (i) abstract the distributed local memory in each core as a shared memory and directly access data placed in a remote core through NoC; and (ii) explicitly partition computation into distributed cores and use message passing to exchange necessary data. We analyze two types of representative systems: LLM runtime or DNN compilers for shared memory architecture such as GPUs, e.g., Ladder [41]; and the SOTA compiler for distributed on-chip memory architectures, e.g., T10 [24] for GraphCore IPU.
Shared-memory system. A shared-memory-based DNN compiler such as Ladder usually assumes a uniform memory access pattern within the underlying memory hierarchy, which cannot tolerate the 1000$\times$ latency variance in wafer-scale accelerators when accessing data from remote memory (failing in L). Moreover, these compilers [10, 27, 47, 49, 52, 36, 41] often focus primarily on partitioning computation, with less emphasis on optimizing data partitioning. This approach can easily lead to significant data duplication and violate the memory constraint requirements (failing in M). Finally, these compilers are unaware of the communication distance of each core, poorly addressing the constraint of routing resources.
Distributed-memory system. The T10 system [24] is designed for AI accelerators with an on-chip crossbar which ensures a constant hop of memory access to other cores on the same chip. T10 handles small local memory and balances communication loads, addressing memory constraints (M) and routing resource limits (R). However, on a PLMR device, it fails to account for varying hop distances (failing in L) and scales to thousands, not millions, of cores (failing in P).
@@ 4 Wafer-Scale LLM Parallelism
We present wafer-scale LLM parallelism, featuring new designs across prefill, decode and KV cache management.
@@ 4.1 Prefill parallelism
The parallelism for LLM prefill must ensure compliance with the PLMR model. Key challenges include: (i) Handling multiple large matrices during prefill, requiring effective dimension partitioning to achieve million-core parallelism (P); (ii) Optimizing GEMM operations, which involve further partitioning and overlapping computation and communication, to minimize long-range communication overhead (L), respect local memory constraints (M), and account for limited routing resources (R); and (iii) Handling matrix transposes, which are costly on a NoC (L) but often required for sequential GEMM operations.
@@
We illustrate this partitioning using self-attention and feedforward, as shown in Figure 3. For this discussion, we define the following annotations: the input activation $A$ and weight $W$ are multi-dimensional tensors during the prefill process. $B$ represents the batch size, $L$ the sequence dimension, $E$ the embedding dimension, $H$ the head dimension, and $F$ the hidden dimension in the feedforward block. As shown by 1 , the partitioning layout of $A$ is represented as $BL_{y}E_{x}$, where the $L$ dimension is partitioned along the $Y$-axis of cores, and the $E$ dimension along the $X$-axis of cores. Similarly, all weight matrices ($W_{Q}$, $W_{K}$, $W_{V}$, $W_{in}$, and $W_{out}$) are partitioned across both dimensions.
Designing PLMR-compliant distributed GEMM. We propose replacing conventional GEMM operators, designed for shared memory architectures, with a newly designed PLMR-compliant distributed GEMM during the prefill phase (as shown in 2 of Figure 3). Unlike TPU and GPU systems that primarily rely on allgather operations for GEMM, PLMR-compliant distributed GEMM algorithms achieve high NoC bandwidth utilization while respecting local memory and routing constraints, ensuring compliance with the L, M, and R properties. This PLMR-compliant distributed GEMM is fully described in Section 5.
@@ 4.2 Decode parallelism
The parallelism strategy for LLM decode must address its memory-bandwidth-intensive nature, presenting several challenges: (i) Decode uses smaller matrices than prefill due to limited input sequences and batch sizes, requiring careful parallelization when dimensions are insufficient for partitioning; (ii) The phase heavily relies on GEMV operations, which are less compute-intensive than GEMM, resulting in short computation phases with limited overlap with communication, making GEMV vulnerable to long-range communication overhead on a NoC (L) and requiring adherence to local memory and routing constraints (M and R); and (iii) Sequential GEMV operations introduce costly matrix transpose on a NoC, risking violation of the L property.
Designing fine-grained replication to enable parallelism at minimal communication cost. When tensor dimensions are insufficient to achieve the high parallelism required for decode, we propose fine-grained replication of tensors in LLMs, specifically replicating the sequence dimension, where the sequence length equals the prompt length during prefill phase and equals 1 during the decode phase. This approach offers two key advantages: (i) It improves parallelism and ensures balanced loads across all cores, and (ii) It avoids additional communication operations such as allreduce. As shown by 1 in Figure4, the $E$ dimension is partitioned along the $y$-axis, and the $L$ dimension is replicated along the $x$-axis, represented as $BE_{y}L^{x}$. Weight matrices $W$ are partitioned across both dimensions, consistent with the prefill phase.
@@
Pre-optimizing model weight placement to avoid matrix transpose. To avoid matrix transpose during decode, we propose pre-optimizing the model weight layout for decode, particularly for the distributed GEMV operation, to eliminate matrix transpose. While this introduces re-placement overhead between prefill and decode phases, the overhead is far smaller than that of sequential matrix transpose during token generation.
Figure 4 illustrates this proposal, detailed in 3 . Specifically, we optimize the placement of weights such as $W_{O}$ and $W_{out}$ for distributed GEMV in decode, differing from their layout in the prefill phase. This approach also removes the need for transpose operations in calculating $Q@K^{T}$ during decode self-attention.
@@ 4.3 Shift-based KV cache management
Existing concatenate-based management causes skewed core utilization. Current KV cache management methods primarily concatenate newly generated KV vectors to the existing cache. While efficient in shared memory architectures, this concatenate operation leads to highly skewed core utilization on PLMR devices, as shown in 1 of Figure 5, where only core in a row is responsible for storing and computing over the newly generated KV vector. After several token generation steps, this only core quickly becomes the bottleneck, as depicted in 2 of Figure 5, causing skewed memory usage and violating the M in PLMR. Moreover, the imbalanced KV cache distribution across cores results in inefficient parallelism, violating the P property.
@@
The upward shifts utilize all NoC links in parallel, maintaining high performance and satisfying the P property. The physical placement of KV cache aligns with logical continuity, adhering to the L property. This method also fully resolves the M violation issue observed in the last row of cores with the concatenate-based approach.
@@ 4.4 Implementation details
Prefill and decode transition. Prefill and decode require distinct strategies. To handle the transition efficiently, we reshuffle KV cache and weights through the fast NoC which often provides 100s Pbits/s aggregated bandwidth, completing instantly without relying on slower off-chip memory.
Parallelism configuration. We empirically determine the scalable parallelism for LLM operators. Automatic parallelism configuration is left for future work.
@@ 5.1 PLMR compliance in distributed GEMM
To identify an scalable distributed GEMM for PLMR devices, we define the following metrics: (i) Paths per core: The number of routing paths per core, with fewer paths ensuring compliance with the R property. (ii) critical path: The longest communication path in each step to transmit submatrix (as the red lines in Figure 6), with fewer hops adhering to the L property. (iii) Memory per core: The memory required per core, with lower usage ensuring the M property.
@@
    GEMM via Allgather is commonly used in GPU and TPU pods for distributed GEMM [34, 51, 31]. Its longest communication path in each step is one core gathering data from the farthest cores, shown as the red line in Figure 6  1 , and $N$ steps to complete the allgather. Each core creates $N$ communication paths to neighbors in its row and column (violating R). The gather in each step spans the critical path with $O(N)$ hops (violating L), and each core uses $O(1/N)$ memory due to inflated working buffers, far exceeding the $O(1/N^{2})$ for local submatrices (violating M).
@@
    SUMMA is Cerebras’ default choice for distributed GEMM on its wafer-scale engine [8]. Its longest communication path in each step is where one core broadcasts data to the farthest core along the column or row, shown by the red line in 2 of Figure 6. Every core creates $N$ communication paths (violating R) and spans the critical path with $O(N)$ hops (violating L) in the longest path. While SUMMA improves memory usage compared to AllGather, requiring only a working set equal to the size of locally partitioned submatrices, it still doubles memory usage.
@@
    Cannon is mesh-optimized choice for distributed GEMM [7], popular in supercomputers. Its longest communication path in each step is the head cores send data to the tail cores. As shown in 3 of Figure 6, each core communicates with two neighbours in a 2D torus, and only needs $O(1)$ communication paths and optimal memory usage of $O(1/N^{2})$. However, it incurs the critical path with $O(N)$ hops as the red line, violating L.
@@
    MeshGEMM (Ours) is a distributed GEMM which complies with the PLMR model. Its longest communication path in each step is shown as the red line in 4 of Figure 6. Each core communicates with two neighbors, two hops away (proven in later sections to be scalable for mesh architectures). This design achieves $O(1)$ communication paths per core needed and optimal memory usage of $O(1/N^{2})$, similar to Cannon. Crucially, it bounds the critical path to 2 hops with $O(1)$ complexity, making it uniquely capable of addressing the L property.
@@ 5.2 Design intuitions and scalability analysis
Our design involves two steps: (i) We ensure algorithm correctness using a cyclic shifting process for GEMM, and (ii) We prove that two-hop communication on this cycle is the minimal distance required to satisfy the L property.
@@
Interleaving. For the flatten communication plan, we would like to minimize the length of the critical path further, thus satisfying the L property. Our key intuition here is to introduce an INTERLEAVE operation to find the mapping relationship from logical to physical, defined in Algorithm 1. As shown by 1 of Figure 7, MeshGEMM first insert core 1 in between core 0 and 4 and core 2 in between core 4 and 3 to form a logical mapping, and then call the INTERLEAVE operation to get the send to and receive from neighbours’ index, resulting in a permutated, equivalent communication plan as shown by 2 in Figure 7. For example, there are 5 cores total (N=5), so physical core 2 (index=2) sends data to physical core 4 (send_index=4) and receives from physical core 0 (recv_index=0).
@@ 5.3 The MeshGEMM algorithm
    Initialization: Consider $C=A\times B$. MeshGEMM will partition $A$ and $B$ into tiles $A_{sub}$ and $B_{sub}$ along two dimensions, forming $N\times N$ tiles, which are distributed across the cores. Each core receives one tile of $A_{sub}$ and one of $B_{sub}$. MeshGEMM will then use INTERLEAVE to initialize the neighbor’s positions for each core.
@@ 5.4 Implementation details
Transposed distributed GEMM. The above algorithm key steps can be applied to the computation of $C=A\times B^{T}$, the dist-GEMM-T in Figure 3 to avoid transposing $B$ on mesh. It does not require alignment before computation and only necessitates $N$ steps two-hop compute-shift for the right matrix $B$ along the Y-axis. After each shift step, each core computes $C_{\text{sub}}=A_{\text{sub}}\times B_{\text{sub}}$, followed by a ReduceAdd of $C_{\text{sub}}$ along the X-axis. After $N$ steps, the final matrix $C$ is obtained.
@@ 6.1 PLMR compliance in distributed GEMV
The completion time of a distributed GEMV is primarily determined by an allreduce operation that aggregates partial results from all selected cores and broadcasts the aggregated results back to all cores. So, we define the number of add-operations (hops) in the longest aggregation path as the critical path in GEMV. Below, we analyze common distributed GEMV implementations in LLM systems and demonstrate that MeshGEMV is the only approach fully compliant with the PLMR model.
@@
    GEMV with pipeline allreduce is commonly used in TPU pod systems [34] and as the default in Cerebras demo [9]. As shown by 1 in Figure 8, it bounds routing resource usage to $O(1)$ per core (meeting R in PLMR). However, its longest aggregation path is from tail to head cores, as shown in the red line, and spans the critical path at $O(N)$, violating the L property.
@@
    GEMV with ring allreduce is commonly used in GPU pod systems, where it is the default configuration. As shown by 2 in Figure 8, it bounds routing resource usage to $O(1)$ (meeting R in PLMR). However, it spans $O(N)$ hops in the critical path, violating the L property.
@@
    GEMV with two-way K-tree allreduce (Ours). As shown by 4 in Figure 8, we build a balanced K-tree to reduce from two-way; its longest aggregation path is from the head or tail core to the tree root core. The critical path is $O(\sqrt[K]{N}K)$ which can address the L. The max number of communication paths at each root core is $O(K)$, and can meet the R limitation by adjusting the K.

### v3

@@ 3.2 Limitations of state-of-the-art approaches
Leveraging the PLMR model, we analyze why existing AI systems fail to fully utilize wafer-scale accelerators. To run an LLM on a wafer-scale accelerator, we generally have two choices: (i) abstract the distributed local memory in each core as a shared memory and directly access data placed in a remote core through NoC; and (ii) explicitly partition computation into distributed cores and use message passing to exchange necessary data. We analyze two types of representative systems: LLM runtime or DNN compilers for shared memory architecture such as GPUs, e.g., Ladder [45]; and the SOTA compiler for distributed on-chip memory architectures, e.g., T10 [25] for GraphCore IPU.
Shared-memory system. A shared-memory-based DNN compiler such as Ladder usually assumes a uniform memory access pattern within the underlying memory hierarchy, which cannot tolerate the thousands of times latency variance in wafer-scale accelerators when accessing data from remote memory (failing in L). Moreover, these compilers [10, 28, 51, 54, 57, 37, 45] often focus primarily on partitioning computation, with less emphasis on optimizing data partitioning. This approach can easily lead to significant data duplication and violate the memory constraint requirements (failing in M). Finally, these compilers are unaware of the communication distance of each core, poorly addressing the constraint of routing resources.
Distributed-memory system. The T10 system [25] is designed for AI accelerators with an on-chip crossbar, which ensures a constant latency of memory access to any other cores on the same chip. T10 handles small local memory and balances communication loads, addressing memory constraints (M) and routing resource limitations (R). However, on a PLMR device, it fails to account for varying memory access latency (failing in L) and scales to thousands, not millions, of cores (failing in P).
@@ 4 Wafer-Scale LLM Parallelism
We present wafer-scale LLM parallelism, featuring new designs across prefill, decode, and KV cache management.
@@ 4.1 Prefill parallelism
The parallelism for LLM prefill must ensure compliance with the PLMR model. Key challenges include: (i) Handling multiple large matrices during prefill, requiring effective dimension partitioning to achieve million-core parallelism (P); (ii) Optimizing GEMM operations, which involve further partitioning and overlapping computation and communication, to minimize long-range communication latency (L), respect local memory constraints (M), and account for limited routing resources (R); and (iii) Handling matrix transposes, which are costly on a mesh NoC (L) but often required for sequential GEMM operations.
@@
[…]
Designing PLMR-compliant distributed GEMM. We propose replacing conventional GEMM operators, which are designed for shared memory architectures, with a newly designed PLMR-compliant distributed GEMM during the prefill phase (as shown in 2 of Figure 3). Unlike TPU and GPU systems that primarily rely on allgather operations for GEMM, PLMR-compliant distributed GEMM algorithms achieve high NoC bandwidth utilization while respecting local memory and routing constraints, ensuring compliance with the L, M, and R properties. This PLMR-compliant distributed GEMM is fully described in Section 5.
@@ 4.2 Decode parallelism
The parallelism strategy for LLM decode must address its memory-bandwidth-intensive nature, presenting several challenges: (i) Decode uses smaller matrices than prefill due to limited input sequences and batch sizes, requiring careful parallelization when dimensions are insufficient for partitioning; (ii) The phase heavily relies on GEMV operations, which are less compute-intensive than GEMM, resulting in short computation phases with limited overlap with communication, making GEMV vulnerable to long-range communication latency on a mesh NoC (L) and requiring adherence to local memory and routing constraints (M and R); and (iii) Sequential GEMV operations introduce costly matrix transpose on a NoC, risking violation of the L property.
Designing fine-grained replication to enable parallelism at minimal communication cost. When tensor dimensions are insufficient to achieve the high parallelism required for decode, we propose fine-grained replication of tensors in LLMs, specifically replicating the sequence dimension, where the sequence length equals the prompt length during prefill phase and equals 1 during the decode phase. This approach offers two key advantages: (i)  it improves parallelism and ensures balanced loads across all cores, and (ii)  it avoids additional communication operations such as allreduce across all cores. As shown by 1 in Figure 4, the $E$ dimension is partitioned along the $y$-axis, and the $L$ dimension is replicated along the $x$-axis, represented as $BE_{y}L^{x}$. Weight matrices $W$ are partitioned across both dimensions, consistent with the prefill phase.
@@
Pre-optimizing model weight placement to avoid matrix transpose. To avoid matrix transpose during decode, we pre-optimize the model weight layout for decode, particularly for the distributed GEMV operation, to eliminate matrix transpose. While this introduces re-placement overhead between prefill and decode phases, the overhead is far smaller than that of sequential matrix transpose during token generation.
Figure 4 illustrates this proposal, detailed in 3 . Specifically, we optimize the placement of weights such as $W_{O}$ and $W_{out}$ for distributed GEMV in decode, differing from their layout in the prefill phase. This approach also removes transpose operations in calculating $Q@K^{T}$ during decode self-attention.
@@ 4.3 Shift-based KV cache management
Existing concatenate-based management causes skewed core utilization. Current KV cache management methods primarily concatenate the most recently generated KV vectors to the existing cache. Though efficient in shared memory architectures, this concatenate operation leads to highly skewed core utilization on PLMR devices, as shown in 1 of Figure 5, where only the core in a row is responsible for storing and computing over the generated KV vector. After several token generation steps, this only core quickly becomes the bottleneck, as depicted in 2 of Figure 5, causing skewed memory usage and violating the M in PLMR. Moreover, the imbalanced KV cache distribution across cores results in inefficient parallelism, violating the P property.
@@
The upward shifts utilize NoC links in parallel, maintaining high performance and satisfying the P property. The physical placement of KV cache aligns with logical continuity and only involves data movement between the adjacent cores, adhering to the L property. This method also fully resolves the M violation issue observed in the last row of cores with the concatenate-based approach.
@@ 4.4 Implementation details
Prefill and decode transition. Prefill and decode require distinct strategies. To handle the transition efficiently, we reshuffle KV cache and weights through the fast NoC, which often provides 100s Pbits/s aggregated bandwidth, completing instantly without relying on slower off-chip memory.
Parallelism configuration. WaferLLM uses offline autotuning to select core counts for each model, optimizing latency based on model size, input/output length, memory per core, and prefill/decode phases. It chooses different core counts for each phase, with fast dynamic remapping enabled by high NoC bandwidth. For models with variable input/output lengths, average values are used to maintain near-peak performance. Autotuning runs separately per model to adapt to specific needs.
@@ 5.1 PLMR compliance in distributed GEMM
To identify a scalable distributed GEMM for PLMR devices, we define the following metrics: (i) Routing paths per core: The number of routing paths per core, with fewer paths ensuring compliance with the R property. (ii) Latency of critical path: Maximal latency among all communication paths in each step to transmit submatrix (as the red lines in Figure 6), with less latency adhering to the L property. (iii) Memory per core: The memory required per core, with lower usage ensuring the M property.
@@
    GEMM via Allgather is commonly used in GPU and TPU pods for distributed GEMM [35, 56, 32]. Its critical communication path in each step is one core gathering data from the farthest cores, shown as the red line in Figure 6  1 , and $N$ steps to complete the allgather. Each core creates $N$ routing paths to neighbors in its row and column (violating R). The constraint of $R$ necessitates step-by-step submatrix transmission via intermediate cores, introducing a communication latency of $O[(\alpha+\beta)N]$ along the critical path (violating L), and each core uses $O(1/N)$ memory due to inflated working buffers (violating M).
@@
    SUMMA is Cerebras’ default choice for distributed GEMM on its wafer-scale engine [7]. Its critical communication path in each step is where one core broadcasts data to the farthest core along the column or row, shown by the red line in 2 of Figure 6. Same to the allgather communication in  1 , each core creates $N$ routing paths (violating R) and spans the critical path with $O[(\alpha+\beta)N]$ latency (violating L) in each step. While SUMMA improves memory usage compared to allgather, requiring only a working set equal to the size of locally partitioned submatrices, it still doubles the peak memory usage.
@@
    Cannon is mesh-optimized choice for distributed GEMM [6], popular in supercomputers and distributed cluster. Its critical communication path in each step is the head cores send data to the tail cores shown in  3 of Figure 6. Each intermediate core communicates with two neighbours in a 2D torus and passes through the submatrix from head to tail, which only needs $O(1)$ communication paths and optimal memory usage of $O(1/N^{2})$.
    Notably, Cannon involves only a constant number of routing paths per core, allowing static routing rules to be assigned for both neighbor communication and the critical path. At each step, submatrices can pass directly from the head to the tail core through intermediate cores, unlike GEMM via allgather or SUMMA, which require step-by-step message transmission. As a result, the critical path incurs only per-hop latency $\alpha$, without additional per-routing overhead $\beta$. However, since the critical path spans $N$ hops, it still incurs $O(\alpha N)$ latency, shown as the red line in 3 , violating the $L$.
@@
    MeshGEMM (Ours) is a distributed GEMM which complies with the PLMR model. The critical communication path in each step, which we named as two-hop transmission, is shown as the red line in 4 of Figure 6, which is significantly shorter than the others. Each core communicates with two two-hop away neighbors(proven in later sections to be scalable for larger-sized mesh architectures), and passes through a one-hop neighbor’s communication. This design achieves $O(1)$ communication paths per core needed and optimal memory usage of $O(1/N^{2})$, similar to Cannon. Crucially, it bounds the critical path to constantly 2 hops with $O(\alpha)$ complexity, making it uniquely capable of addressing the L property.
@@ 5.2 Design intuitions and scalability analysis
Our design involves two phases: (i) We ensure algorithm correctness using a cyclic shifting process for GEMM, and (ii)  then use interleaving to bound the critical path latency to a constant. Based on the cyclic shifting and interleaving, we get the two-hop transmission communication path, and we prove that it, on this cycle, is the minimal distance required to satisfy the L property.
@@
Interleaving. For the flattened communication plan, we would like to minimize the length of the critical path further, thus satisfying the L property. Our key intuition here is to introduce an INTERLEAVE operation to find the mapping relationship from logical to physical, defined in Algorithm 1. As shown by 1 of Figure 7, MeshGEMM first insert core 1 in between core 0 and 4 and core 2 in between core 4 and 3 to form a logical mapping, and then call the INTERLEAVE operation to get the send to and receive from neighbours’ index, resulting in a permutated, equivalent communication plan as shown by 2 in Figure 7. For example, there are 5 cores total (N=5), so physical core 2 (index=2) sends data to physical core 4 (send_index=4) and receives from physical core 0 (recv_index=0). Figures 6 and 7 illustrate the case with 5 cores as an example, while Algorithm 1 demonstrates that the two-hop communication pattern in MeshGEMM generalizes to mesh architectures of arbitrary size with $N\geq 3$.
@@ 5.3 The MeshGEMM algorithm
    Initialization: Consider $C=A\times B$. MeshGEMM will partition $A$ and $B$ into tiles $A_{sub}$ and $B_{sub}$ along two dimensions, forming $N\times N$ tiles, which are distributed across the cores. Each core receives one tile of $A_{sub}$ and one of $B_{sub}$. MeshGEMM will then use INTERLEAVE to initialize the neighbors’ positions for each core.
@@ 5.4 Implementation details
Transposed distributed GEMM. The above algorithm can be applied to the computation of $C=A\times B^{T}$, the dist-GEMM-T in Figure 3 to avoid transposing $B$ on mesh. It does not require alignment before computation and only necessitates $N$ steps of two-hop compute-shift for the right matrix $B$ along the Y-axis. After each shift step, each core computes $C_{\text{sub}}=A_{\text{sub}}\times B_{\text{sub}}$, followed by a ReduceAdd of $C_{\text{sub}}$ along the X-axis. After $N$ steps, the final matrix $C$ is obtained.
@@ 6.1 PLMR compliance in distributed GEMV
The completion time of a distributed GEMV is primarily determined by an element-wise computation to generate a partial sum result at each core and then an allreduce operation that aggregates partial results from all selected cores and then sends the aggregated results back to all cores for the continuous GEMV. As the analysis in GEMM, in wafer-scale GEMV, we also define the following metrics: (i) Routing paths per core: The number of routing paths per core, with fewer paths ensuring compliance with the R property (ii) Latency of critical path: Maximal latency of the whole allreduce communication, mainly determined by the time from when the partial sum is sent from the farthest core to when the aggregated sum is received by the farthest core.
We analyze current distributed GEMV methods and show how MeshGEMV meets these metrics:
@@
    GEMV with pipeline allreduce is commonly used in TPU pod systems [35] and as the default in Cerebras demo [9]. As illustrated by the red line in the Figure 8  1 , the critical path of the pipelined all-reduce starts from the farthest core, where partial sums are reduced step-by-step toward the root core, and the aggregated result is then broadcast from root to all cores. While this approach ensures that the routing path per core remains within the $R$ constraint, $O(1)$ per-core routing resource usage, it incurs a total of $2N$ hops and $N$ routing stages, violating the L property.
@@
    GEMV with ring allreduce is commonly used in GPU pod systems, where it is the default configuration, especially for a large amount of data [27]. As shown by 2 in Figure 8, the critical path of ring allreduce involves each partial sum traversing all cores in the ring. Similar to pipelined allreduce, this approach maintains $O(1)$ routing paths per core, but incurs a communication latency of $O\left[(2\alpha+\beta)N\right]$, thus violating the latency constraint $L$ defined in PLMR.
@@
[…]